OpenStudy (anonymous):

write the equation of a line that passes through the point (-6,1) and is parallel to the line 2x-y=7

OpenStudy (anonymous):

For the lines to be parallel, the gradients must be the same, so the left hand side of the equation doesn't need changing. Now all you have to do is calculate the constant on the right:\[2 \times -6 -(-1)=-11\]giving an answer of\[y-2x=11\]

OpenStudy (anonymous):

Thanks!